how fast must a car tavel to overtake a bus in 5 hours, if the bus is averaging 30 kilometers per hour and left 3 hours before the car
Answer by Gogonati(855)

Solution:
Denote x km/h the speed of the car. For 5 hours its traveling distance is
5*x km while the bus for 5 hours travel the distance: 5*30=150 km,since the bus
left 3 hours before the car its total distance from the car is 150+3*30=240 km.
The car overtake the bus when 5*x=240 where x=48 km/h.
Answer:The car must move 48 km/h.